40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Entry  Wed Feb 8 17:01:01 2017, Gabriele, Optics, Daily Progress, Two laser installed, setup aligned photo_setup_new.png
    Reply  Thu Feb 9 09:30:52 2017, Gabriele, Optics, Daily Progress, Two laser installed, setup aligned Screenshot.png
Message ID: 300     Entry time: Thu Feb 9 09:30:52 2017     In reply to: 296
Author: Gabriele 
Type: Optics 
Category: Daily Progress 
Subject: Two laser installed, setup aligned 

The lasers are behaving well, there is no high noise or wandering lines. The spectrum below is taken in air: that explains the excess of noise in the few kHz region.

Quote:

Today I swapped out the 8mW laser and installed two 2.5 mW lasers. I rebuilt the input part of the optical levers and re-aligned everything. See below for a picture of the new setup: red beams are input, yellow beams are output. I also installed a protective screen all around the table, to abvoid any suprios beam to get out.

 

ELOG V3.1.3-